Assume that your game piece is on North Carolina Avenue and that there are hotels (that are not yours!) on Park Place and Boardwalk. You roll a pair of dice on your next turn. Hint: North Carolina Avenue, Park Place, and Boardwalk are all on the right side of the game board below.

a. What are the odds that you land on one of these properties?

b. What is the probability that you avoid these properties on your next move? Write your answer as a fraction.
Answer #1

ANSWER:

Game piece is on North Carolina Avenue and that there are hotels (that are not yours!) on Park Place and Boardwalk. You roll a pair of dice on your next turn. Hint: North Carolina Avenue, Park Place, and Boardwalk are all on the right side of the game board below. a. What are the odds that you land on one of these properties? b. What is the probability that you avoid these properties on your next move? Write your answer as a fraction.

MONOPOLY game with 2 dice each having 6 numbers: 1,2,..6.Thus, total number of all possibilities= 6 X 6= 36. The different possibilities are (1,1), (1,2),...,(6,6). They are 36 possibilities in total. MONOPOLY game is ALWAYS played with 2 dice each marked with 6 NUMBERS from 1 to 6. This is a pre-requite of playing MONOPOLY.

(a)

Odds that you land on one of these two propoerties (i) Parkplace and (ii) Boardwalk = Number of ways you land on these two properties/ Number of ways you do not land on these two properties.

(i) Park Place you land by putting 5

Number of ways of putting 5: (1,4), (2,3), (3,2), (4,1): 4 Ways

(ii) Boardwalk you land by putting 7

Number of ways of putting 7: (1,6), (2,5), (3,4), (4,3), (5,2), (6,1) : 6 ways

Number of ways you land on these two properties = 10 Ways

Number of ways you do not land on these properties = 36 - 10 = 26 Ways

Substituting, we get:

Odds that you land on one of these properties = 10:26

                                                                    = 5 : 13

(b)

Probability that you avoid these two properties = 26/36

                                                              = 13/18